# Optimal Switching in Networked Control Systems: Finite Horizon

### 29 Mar 2026 v1

@@ 7 Numerical Results
In this section, we evaluate the performance of the optimal policy (44),(82) and compare it against commonly used sub-optimal controller and switching-policy combinations. As established earlier, the optimal policy applies to any zero-mean disturbance distribution symmetric about the origin and does not rely on Gaussian assumptions. We therefore evaluate its performance under multiple symmetric noise models.
@@
In contrast, the impulsive controller concentrates the entire control action at the observation instant and applies no control between switching times:
@@
Since the impulsive controller exerts no input during the inter-switching interval, the feedback gain is chosen as a/b-a/b rather than L-L to ensure a fair comparison. Note that this gain corresponds to the limiting case of the LQR controller as $r\to 0$, where control effort is penalized negligibly and the state is driven to zero in a single step.
As benchmark switching strategies, we consider three alternatives: random (Bernoulli), periodic, and symmetric threshold (SYM) policies.
The random (Bernoulli) switching policy sets $\{D_{k}\}$ as an i.i.d. Bernoulli sequence with
@@
The periodic policy is defined in Theorem 3. It is the “as periodic as possible” policy that satisfies the rate constraint for any $0<r_{s}<1$ by distributing switching instants as uniformly as possible over time.
Figure 3: Running-average LQR cost versus time under Gaussian disturbances. All policies are tuned to satisfy the same target switching rate $r_{s}=0.4$. The optimal symmetric threshold policy (OPT) achieves the lowest cost, followed closely by symmetric policies, while periodic and random strategies incur higher cost. Results are averaged over 100 Monte Carlo runs.
Figure 5: Steady-state LQR cost versus the target switching rate $r_{s}$ for $a=1$ under Gaussian disturbances. ZOH policies become unstable for $r_{s}\lesssim 0.4$, while the optimal policy achieves the lowest cost across all rates. As $r_{s}\to 1$, all policies converge to similar performance.
As the last switching policy, we consider the symmetric switching policy defined in (82), where the threshold is chosen so that the resulting switching rate matches $r_{s}$. Although the symmetric switching policy is optimal when paired with the optimal controller, it becomes suboptimal when combined with ZOH or impulsive controllers, since the controller itself is no longer optimal. Note that in the limiting case $r\to 0$, the optimal controller (9) converges to the impulsive controller (84). Consequently, as $r\to 0$, the optimal policy (44)–(82) coincides with the impulsive-controller and symmetric-switching combination.
All of the sub-optimal switching policies defined above satisfy the symmetry condition in (24). To broaden the range of policies under consideration, we also introduce a non-symmetric switching rule, referred to as the state-based policy, defined as
@@
since the switching decision now depends on the most recent state observation $X_{t_{k}}$, thereby breaking the symmetry.
A key consequence is that the controller no longer retains the simple structure in (44). Instead, the conditional state estimate $\hat{X}_{k}$ incorporates bias terms that depend explicitly on the noise distribution and the switching decisions, as reflected in the modified estimation error dynamics (42). More precisely, the optimal controller takes the form
@@
We consider a scalar LTI system with parameters $a=1$, $b=1$, $q=1$, $r=1$, disturbance variance $\sigma_{W}=10$, and delay $\tau=1$, unless stated otherwise. The communication constraint is imposed through a maximum allowable switching rate, denoted by $r_{s}$. Since the switching rate induced by threshold-based policies depends implicitly on the chosen threshold, we tune the thresholds of all policies so that their empirical switching rates match $r_{s}$ as closely as possible. This ensures a fair comparison across different policies under a common communication budget.
@@
We next examine how the relative performance of these policies varies with the open-loop gain $a$, the switching rate $r_{s}$, and the disturbance variance $\sigma_{W}$.
@@
In Fig. 5, we plot the steady-state LQR cost as a function of the target switching rate $r_{s}$ for $a=1$ under Gaussian disturbances. The optimal policy consistently achieves the lowest cost across all switching rates. As the communication constraint becomes more stringent (i.e., as $r_{s}$ decreases), ZOH policies rapidly degrade and become unstable. In particular, all ZOH policies diverge for $r_{s}<0.4$, consistent with the observations in Fig. 3. While the confidence intervals appear larger in this figure, this is primarily due to the scaling of the vertical axis. Finally, as $r_{s}\to 1$, all policies exhibit similar performance, reflecting the diminishing impact of communication constraints when transmissions are nearly always allowed.
@@
In Fig. 7, we plot the steady-state LQR cost as a function of the target switching rate $r_{s}$ for $a=1$ under Laplace disturbances. Comparing with Fig. 5, we observe qualitatively similar behavior across all policies. In particular, ZOH-based policies become unstable for $r_{s}<0.4$, and the state-based policy consistently provides the second-best performance for moderate to large switching rates ($r_{s}>0.3$).
@@
In Fig. 8, we plot the normalized difference in steady-state LQR cost between non-Gaussian and Gaussian disturbances as a function of the open-loop gain $a$ for the optimal policy. The results show that the noise distribution has a noticeable impact on performance. In particular, the gap between Laplace and Gaussian costs decreases as $a$ increases, indicating that the effect of the distribution diminishes for more unstable systems. In contrast, Fig. 9 shows the same comparison for the periodic impulsive policy. Here, the performance under different noise distributions remains nearly identical across all values of $a$. This can be attributed to the fact that, under periodic switching, the transmission decisions are independent of the system state and noise realizations. As a result, the interaction between the noise distribution and the switching mechanism is eliminated, leading to distribution-insensitive performance.
@@
Figure 9: Normalized steady-state cost difference relative to Gaussian disturbances versus $a$ for the periodic impulsive policy. The performance is largely insensitive to the noise distribution due to the independence of switching decisions from system dynamics.
In Fig. 10, we plot the normalized steady-state cost difference relative to the Gaussian case as a function of $r_{s}$ for the optimal policy. As in the previous results, the performance depends on the underlying noise distribution, indicating that the interaction between the switching decisions and the disturbance remains significant. In contrast, Fig. 11 shows that for the periodic impulsive policy, the normalized cost difference is nearly identical for both Laplace and uniform disturbances across all values of $r_{s}$. This suggests that the performance of such policies is largely insensitive to the specific noise distribution.
This behavior can be explained by the fact that periodic (self-triggered) switching policies are independent of the noise realizations. Consequently, for any zero-mean symmetric disturbance distributions with identical variance, the induced closed-loop cost is effectively invariant to the choice of distribution.
Figure 10: Normalized steady-state cost difference relative to Gaussian disturbances versus $r_{s}$ for the optimal policy. The performance varies with the noise distribution across all switching rates.
@@ 8 Conclusion
This paper studied the constrained joint design of switching and control policies for scalar networked control systems under an average switching-rate constraint. The main difficulty arises from the dual effect: switching decisions influence the controller’s information pattern and hence the estimation process, which in general prevents the direct application of the classical separation principle. To address this issue, we carefully analyzed the underlying information structure and identified the class of symmetric policies, under which the accumulated disturbance remains conditionally zero-mean at the controller.
Our main result established that this class is not merely a tractable restriction, but is in fact optimal for the finite-horizon switched LQR problem. As a consequence, separation holds under the optimal policy: the optimal controller retains the standard linear feedback form through the conditional state estimate, while the optimal switching mechanism is characterized by a symmetric threshold structure. In particular, the switching decision can be expressed in terms of the accumulated disturbance, or equivalently the estimation error under symmetry.
@@ Appendix A Proof of Theorem 1
And in Step (c), we expanded the conditional expectation $\mathbb{E}[X_{N-1}^{2}\mid I^{\mathcal{C}}_{N-1}]$. Note that $u_{N-1}^{\star}$ includes all possible switching policies up to N1N-1, i.e. $\boldsymbol{\pi}=(f_{0},f_{1},\dots,f_{N-1})\in\Pi$.
@@
Note that the controller and cost-to-go functions are both dependent on the switching policy. Since we have not found the optimal switching policy, the controller is not fully determined, either. The analysis shows that the optimal controller is a linear function of the estimator. Hence, the effect of the delay can only be seen through $\mathbb{E}[X_{k}\mid I^{\mathcal{C}}_{k}]$ in $u_{k}^{\star}$.
@@ Appendix B Proof of Theorem 2
The proof follows mostly from [5, Chapter 8, Lemma 6.1]. However, noisy observations have been considered in [5] without a switch and delay. We will show that the proof is valid for both intermittent observations and delay.
@@
$\overset{(b)}{\approx}\frac{1}{N}\left(P_{0}\,\mathbb{E}[X_{0}^{2}]+\sum_{k=0}^{N-1}L\left(r+b^{2}P\right)\left(X_{k}-\mathbb{E}[X_{k}|I^{\mathcal{C}}_{k}]\right)^{2}+NP\,\sigma_{W}^{2}\right)$
$=\frac{P_{0}}{N}\,\mathbb{E}[X_{0}^{2}]+\frac{1}{N}\sum_{k=0}^{N-1}L\left(r+b^{2}P\right)\left(X_{k}-\mathbb{E}[X_{k}|I^{\mathcal{C}}_{k}]\right)^{2}+P\,\sigma_{W}^{2},$ (100)
@@ Appendix C Proof of Theorem 3
Assume that the initial budget is $Q_{0}=\lfloor N\,r_{s}\rfloor$ and the system start at $k=-\tau$ with a switching $t_{0}=-\tau$. Also assume that the last switching occurs at $t_{Q_{0}+1}=N-\tau$. Note that the last effective decision occurs at $N-\tau-1$, hence this assumption does not break causality. Hence, we can rewrite the cost in $P_{2}$ as
@@
where $\Delta_{n}=t_{n+1}-t_{n}$ is the inter-switching time.
@@
where the minimization is taken over admissible switching policies $\pi$. Where the constraint $\sum_{k=0}^{N-1}\Delta_{k}=\lfloor Nr_{s}\rfloor=Q_{0}$ follows from the fixed switching-rate requirement. Over a horizon of length $N$, a switching rate $r_{s}$ permits at most $Nr_{s}$ switch activations. Since switching times are integer-valued, the total number of switches must equal $\lfloor Nr_{s}\rfloor$, yielding the stated constraint.
@@
In both cases, the objective depends on the inter-switching times only through a separable convex function of $\Delta_{n}$. Indeed, for $a^{2}\neq 1$ the nonconstant term $\mathbb{E}[a^{2\Delta_{n}}]$ is convex in $\Delta_{n}$, while for $a^{2}=1$ the term $\mathbb{E}[\Delta_{n}^{2}]$ is strictly convex. Consequently, both problems reduce to an integer convex partitioning problem of the form
@@
By convexity of the cost, transferring one unit of time from a larger interval to a smaller one preserves feasibility and does not increase the objective, with strict improvement when the cost is strictly convex. Repeating this balancing step eliminates such disparities and yields an optimal solution in which all inter-switching times differ by at most one, i.e.,
@@
occurrences of $\left\lfloor N/Q_{0}\right\rfloor+1$. Placing these longer intervals as evenly as possible over the horizon yields an “as periodic as possible” switching policy, which is therefore optimal for both (114) and (115).
@@ Appendix D Proof of Lemma 2
Here, step (a) follows from the DP recursion (34) and the fact that, at stage $k=\ell$ with budget $Q_{k}=\bar{s}$, the switch can transmit at all subsequent time instants. Consequently, the minimizing action is $D_{k}=1$, and only the corresponding stage cost is incurred.
Similarly, to establish (41), observe that at stage $\ell$, if the switch chooses not to transmit, i.e., $D_{\ell}=0$, and the remaining budget satisfies $Q_{\ell}=s-1$, then switching can be performed at all subsequent time instants. Consequently, (41) follows directly from the DP recursion (34) and the next-stage value function under switching given by (40).
@@ Appendix E Proof of Theorem 4
$\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\mathbb{E}\!\left[V^{\prime}_{\ell(s+1)0}(I_{\ell}^{\mathcal{S}})\right].$ (117)
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\right]$ $=\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\qquad\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\mid D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]$ (118)
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right].$
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]$
$\qquad=\mathbb{E}\!\left[a^{2\tau}\bar{\mathcal{E}}_{\ell}^{\,2}+s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)\,\middle|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]$
$\qquad\overset{(a)}{=}\mathbb{E}\!\Bigg[a^{2\tau}\!\left(S_{m}-\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m}=1,\;D_{\ell-m+1}=0,\;\cdots,D_{\ell}=0\right]\right)^{\!2}$ (119)
$\qquad\qquad+s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)\,\Bigg|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Bigg].$ (120)
@@
$=\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right]$
$=\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)\;\mathbb{E}\!\Bigg[s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)$
$\qquad+\;a^{2\tau}\left(S_{m}-\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m}=1,\;D_{\ell-m+1}=\cdots=D_{\ell}=0\right]\right)^{\!2}$
$\qquad\qquad\qquad\,\Bigg|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Bigg]$
$\overset{(a)}{=}a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\qquad\mathbb{E}\!\Bigg[\left(S_{m}-C_{m}(X_{\ell-m},\pi)\right)^{\!2}\Bigg|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Bigg].$ (121)
@@
$C_{m}(X_{\ell-m},\pi)\triangleq\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m}=1,\;D_{\ell-m+1}=\cdots=D_{\ell}=0\right]$ (122)
@@
$\mathbb{E}\!\Big[(S_{m}-C_{m}(X_{\ell-m},\pi))^{2}\,\Big|\,D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad=\int_{-\infty}^{\infty}\mathbb{E}\!\Big[(S_{m}-C_{m}(x,\pi))^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m}=1,\;Q_{\ell-m}=s-1\Big]\,f^{\prime}_{X}(x)\,dx.$ (123)

### 02 May 2026 v3

@@ 7 Numerical Results
In this section, we evaluate the performance of the optimal policy (44),(82) and compare it against commonly used sub-optimal controller and scheduling policy combinations. As established earlier, the optimal policy applies to any zero-mean disturbance distribution symmetric about the origin and does not rely on Gaussian assumptions. We therefore evaluate its performance under multiple symmetric noise models.
@@
In contrast, the impulsive controller concentrates the entire control action at the observation instant and applies no control between update times:
@@
Since the impulsive controller exerts no input during the inter-scheduling interval, the feedback gain is chosen as a/b-a/b rather than L-L to ensure a fair comparison. Note that this gain corresponds to the limiting case of the LQR controller as $r\to 0$, where control effort is penalized negligibly and the state is driven to zero in a single step.
Figure 5: Steady-state LQR cost versus the target communication rate $r_{s}$ for $a=1$ under Gaussian disturbances. ZOH policies become unstable for $r_{s}\lesssim 0.4$, while the optimal policy achieves the lowest cost across all rates. As $r_{s}\to 1$, all policies converge to similar performance.
As benchmark scheduling strategies, we consider three alternatives: random (Bernoulli), periodic, and symmetric threshold (SYM) policies.
The random (Bernoulli) scheduling policy sets $\{D_{k}\}$ as an i.i.d. Bernoulli sequence with
@@
The periodic policy is defined in Theorem 3. It is the “as periodic as possible” policy that satisfies the rate constraint for any $0<r_{s}<1$ by distributing scheduling instants as uniformly as possible over time.
As the last scheduling policy, we consider the symmetric scheduling policy defined in (82), where the threshold is chosen so that the resulting communication rate matches $r_{s}$. Although the symmetric scheduling policy is optimal when paired with the optimal controller, it becomes suboptimal when combined with ZOH or impulsive controllers, since the controller itself is no longer optimal. Note that in the limiting case $r\to 0$, the optimal controller (9) converges to the impulsive controller (84). Consequently, as $r\to 0$, the optimal policy (44),(82) coincides with the impulsive-controller and symmetric-scheduling combination.
All of the sub-optimal scheduling policies defined above satisfy the symmetry condition in (24). To broaden the range of policies under consideration, we also introduce a non-symmetric scheduling rule, referred to as the state-based policy, defined as
@@
since the scheduling decision now depends on the most recent state observation $X_{t_{k}}$, thereby breaking the symmetry.
A key consequence is that the controller no longer retains the simple structure in (44). Instead, the conditional state estimate $\hat{X}_{k}$ incorporates bias terms that depend explicitly on the noise distribution and the scheduling decisions, as reflected in the modified estimation error dynamics (42). More precisely, the optimal controller takes the form
@@
We consider a scalar LTI system with parameters $a=1$, $b=1$, $q=1$, $r=1$, disturbance variance $\sigma_{W}=10$, and delay $\tau=1$, unless stated otherwise. The communication constraint is imposed through a maximum allowable communication rate, denoted by $r_{s}$. Since the communication rate induced by threshold-based policies depends implicitly on the chosen threshold, we tune the thresholds of all policies so that their empirical communication rates match $r_{s}$ as closely as possible. This ensures a fair comparison across different policies under a common communication budget.
@@
We next examine how the relative performance of these policies varies with the open-loop gain $a$, the communication rate $r_{s}$, and the disturbance variance $\sigma_{W}$.
Figure 9: Normalized steady-state cost difference relative to Gaussian disturbances versus $a$ for the periodic impulsive policy. The performance is largely insensitive to the noise distribution due to the independence of scheduling decisions from system dynamics.
@@
In Fig. 5, we plot the steady-state LQR cost as a function of the target communication rate $r_{s}$ for $a=1$ under Gaussian disturbances. The optimal policy consistently achieves the lowest cost across all communication rates. As the communication constraint becomes more stringent (i.e., as $r_{s}$ decreases), ZOH policies rapidly degrade and become unstable. In particular, all ZOH policies diverge for $r_{s}<0.4$, consistent with the observations in Fig. 3. While the confidence intervals appear larger in this figure, this is primarily due to the scaling of the vertical axis. Finally, as $r_{s}\to 1$, all policies exhibit similar performance, reflecting the diminishing impact of communication constraints when transmissions are nearly always allowed.
@@
In Fig. 7, we plot the steady-state LQR cost as a function of the target communication rate $r_{s}$ for $a=1$ under Laplace disturbances. Comparing with Fig. 5, we observe qualitatively similar behavior across all policies. In particular, ZOH-based policies become unstable for $r_{s}<0.4$, and the state-based policy consistently provides the second-best performance for moderate to large communication rates ($r_{s}>0.3$).
@@
In Fig. 8, we plot the normalized difference in steady-state LQR cost between non-Gaussian and Gaussian disturbances as a function of the open-loop gain $a$ for the optimal policy. The results show that the noise distribution has a noticeable impact on performance. In particular, the gap between Laplace and Gaussian costs decreases as $a$ increases, indicating that the effect of the distribution diminishes for more unstable systems. In contrast, Fig. 9 shows the same comparison for the periodic impulsive policy. Here, the performance under different noise distributions remains nearly identical across all values of $a$. This can be attributed to the fact that, under periodic scheduling, the transmission decisions are independent of the system state and noise realizations. As a result, the interaction between the noise distribution and the scheduling mechanism is eliminated, leading to distribution-insensitive performance.
@@
In Fig. 10, we plot the normalized steady-state cost difference relative to the Gaussian case as a function of $r_{s}$ for the optimal policy. As in the previous results, the performance depends on the underlying noise distribution, indicating that the interaction between the scheduling decisions and the disturbance remains significant. In contrast, Fig. 11 shows that for the periodic impulsive policy, the normalized cost difference is nearly identical for both Laplace and uniform disturbances across all values of $r_{s}$. This suggests that the performance of such policies is largely insensitive to the specific noise distribution.
This behavior can be explained by the fact that periodic (self-triggered) scheduling policies are independent of the noise realizations. Consequently, for any zero-mean symmetric disturbance distributions with identical variance, the induced closed-loop cost is effectively invariant to the choice of distribution.
Figure 10: Normalized steady-state cost difference relative to Gaussian disturbances versus $r_{s}$ for the optimal policy. The performance varies with the noise distribution across all communication rates.
@@ 8 Conclusion
This paper studied the constrained joint design of scheduling and control policies for scalar networked control systems under an average communication-rate constraint. The main difficulty arises from the dual effect: scheduling decisions influence the controller’s information pattern and hence the estimation process, which in general prevents the direct application of the classical separation principle. To address this issue, we carefully analyzed the underlying information structure and identified the class of symmetric policies, under which the accumulated disturbance remains conditionally zero-mean at the controller.
Our main result established that this class is not merely a tractable restriction, but is in fact optimal for the finite-horizon communication-constrained LQR problem. As a consequence, separation holds under the optimal policy: the optimal controller retains the standard linear feedback form through the conditional state estimate, while the optimal scheduling mechanism is characterized by a symmetric threshold structure. In particular, the scheduling decision can be expressed in terms of the accumulated disturbance, or equivalently the estimation error under symmetry.
@@ Appendix A Proof of Theorem 1
And in Step (c), we expanded the conditional expectation $\mathbb{E}[X_{N-1}^{2}\mid I^{\mathcal{C}}_{N-1}]$. Note that $u_{N-1}^{\star}$ includes all possible scheduling policies up to N1N-1, i.e. $\boldsymbol{\pi}=(f_{0},f_{1},\dots,f_{N-1})\in\Pi$.
@@
Note that the controller and cost-to-go functions are both dependent on the scheduling policy. Since we have not found the optimal scheduling policy, the controller is not fully determined, either. The analysis shows that the optimal controller is a linear function of the estimator. Hence, the effect of the delay can only be seen through $\mathbb{E}[X_{k}\mid I^{\mathcal{C}}_{k}]$ in $u_{k}^{\star}$.
@@ Appendix B Proof of Theorem 2
The proof follows mostly from [5, Chapter 8, Lemma 6.1]. However, noisy observations have been considered in [5] without a scheduling mechanism and delay. We will show that the proof is valid for both intermittent observations and delay.
@@
$\overset{(b)}{\approx}\frac{1}{N}\left(P_{0}\,\mathbb{E}[X_{0}^{2}]+\mathbb{E}\left[\sum_{k=0}^{N-1}L\left(r+b^{2}P\right)\left(X_{k}-\mathbb{E}[X_{k}|I^{\mathcal{C}}_{k}]\right)^{2}\right]+NP\,\sigma_{W}^{2}\right)$
$=\frac{P_{0}}{N}\,\mathbb{E}[X_{0}^{2}]+\frac{1}{N}\mathbb{E}\left[\sum_{k=0}^{N-1}L\left(r+b^{2}P\right)\left(X_{k}-\mathbb{E}[X_{k}|I^{\mathcal{C}}_{k}]\right)^{2}\right]+P\,\sigma_{W}^{2},$ (100)
@@ Appendix C Proof of Theorem 3
Assume that the initial budget is $Q_{0}=\lfloor N\,r_{s}\rfloor$ and the system start at $k=-\tau$ with a scheduling $t_{0}=-\tau$. Also assume that the last scheduling occurs at $t_{Q_{0}+1}=N-\tau$. Note that the last effective decision occurs at $N-\tau-1$, hence this assumption does not break causality. Hence, we can rewrite the cost in $P_{2}$ as
@@
where $\Delta_{n}=t_{n+1}-t_{n}$ is the inter-update time.
@@
where the minimization is taken over admissible scheduling policies $\pi$. Where the constraint $\sum_{k=0}^{N-1}\Delta_{k}=\lfloor Nr_{s}\rfloor=Q_{0}$ follows from the fixed communication-rate requirement. Over a horizon of length $N$, a communication rate $r_{s}$ permits at most $Nr_{s}$ scheduler activations. Since update times are integer-valued, the total number of updates must equal $\lfloor Nr_{s}\rfloor$, yielding the stated constraint.
@@
In both cases, the objective depends on the inter-update times only through a separable convex function of $\Delta_{n}$. Indeed, for $a^{2}\neq 1$ the nonconstant term $\mathbb{E}[a^{2\Delta_{n}}]$ is convex in $\Delta_{n}$, while for $a^{2}=1$ the term $\mathbb{E}[\Delta_{n}^{2}]$ is strictly convex. Consequently, both problems reduce to an integer convex partitioning problem of the form
@@
By convexity of the cost, transferring one unit of time from a larger interval to a smaller one preserves feasibility and does not increase the objective, with strict improvement when the cost is strictly convex. Repeating this balancing step eliminates such disparities and yields an optimal solution in which all inter-update times differ by at most one, i.e.,
@@
occurrences of $\left\lfloor N/Q_{0}\right\rfloor+1$. Placing these longer intervals as evenly as possible over the horizon yields an “as periodic as possible” scheduling policy, which is therefore optimal for both (114) and (115).
@@ Appendix D Proof of Lemma 2
Here, step (a) follows from the DP recursion (34) and the fact that, at stage $k=\ell$ with budget $Q_{k}=\bar{s}$, the scheduler can transmit at all subsequent time instants. Consequently, the minimizing action is $D_{k}=1$, and only the corresponding stage cost is incurred.
Similarly, to establish (41), observe that at stage $\ell$, if the scheduler chooses not to transmit, i.e., $D_{\ell}=0$, and the remaining budget satisfies $Q_{\ell}=s-1$, then scheduling can be performed at all subsequent time instants. Consequently, (41) follows directly from the DP recursion (34) and the next-stage value function under scheduling given by (40).
@@ Appendix E Proof of Theorem 4
$\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\right].$ (117)
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\right]$ $=\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\qquad\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\mid D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]$ (118)
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right].$
@@
$\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]$
$\qquad=\mathbb{E}\!\left[a^{2\tau}\bar{\mathcal{E}}_{\ell}^{\,2}+s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)\,\middle|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]$
$\qquad\overset{(a)}{=}\mathbb{E}\!\Bigg[a^{2\tau}\!\left(S_{m}-\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m-1}=1,\;D_{\ell-m}=0,\;\cdots,D_{\ell}=0\right]\right)^{\!2}$ (119)
$\qquad\qquad+s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)\,\Bigg|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Bigg].$ (120)
@@
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\Bigg\{\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)\mathbb{E}\!\left[V^{\prime}_{\ell(s-1)0}(I_{\ell}^{\mathcal{S}})\,\middle|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right]\Bigg\}$
$=\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\Bigg\{\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)\;\mathbb{E}\!\Bigg[s\!\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)$
$\qquad+\;a^{2\tau}\left(S_{m}-\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m-1}=1,\;D_{\ell-m}=\cdots=D_{\ell}=0\right]\right)^{\!2}$
$\qquad\qquad\qquad\,\Bigg|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Bigg]\Bigg\}$ (121)
$\overset{(a)}{=}\min_{f_{\ell}(I_{\ell}^{\mathcal{S}})}\;\Bigg\{a^{2\tau}\sum_{m=1}^{N-Q_{0}-\tau-1}\Pr\!\left(D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\right)$
$\qquad\qquad\mathbb{E}\!\Bigg[\left(S_{m}-C_{m}(X_{\ell-m},\pi)\right)^{\!2}\Bigg|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Bigg]\Bigg\}.$ (122)
@@
$C_{m}(X_{\ell-m},\pi)\triangleq\mathbb{E}\!\left[S_{m}\,\middle|\,X_{\ell-m},\;D_{\ell-m-1}=1,\;D_{\ell-m}=\cdots=D_{\ell}=0\right]$ (123)
@@
$\mathbb{E}\!\Big[(S_{m}-C_{m}(X_{\ell-m},\pi))^{2}\,\Big|\,D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big]$
$\qquad=\int_{-\infty}^{\infty}\mathbb{E}\!\Big[(S_{m}-C_{m}(x,\pi))^{2}\,\Big|\,X_{\ell-m}=x,\;D_{\ell-m-1}=1,\;Q_{\ell-m}=s-1\Big]\,f^{\prime}_{X}(x)\,dx.$ (124)